A right triangular prism has volume equal to 288 cm^3. The height of the prism is 3 cm. One of the bases of the triangular face (not the hypotenuse) is equal to 12 cm, determine the length of the hypotenuse (in cm) of the triangular face.

Grade:11

2 Answers

Akash Kumar Dutta
11 years ago

Dear Vivek,

i have seen your previous postings...
among them this is the most easy question.
VOLUME=(base area).(height of prism)=288
                (base area)=288/3=96cm
but base area =1/2.(base).(height of trinangle)=96
give base=12cm => height of triangle= 96.2/12=16cm.
hypotenuse= root( 12^2 + 16^2 ) =  20cm (ANS)

Regards.

Abhishekh kumar sharma
10 years ago

VOLUME=(base area).(height of prism)=288
                (base area)=288/3=96cm
but base area =1/2.(base).(height of trinangle)=96
give base=12cm => height of triangle= 96.2/12=16cm.
hypotenuse= root( 12^2 + 16^2 ) =  20cm (ANS)